Circular No. 6950

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998co IN NGC 7131 Wayne Johnson, Anza Observatory, reports his discovery of a super-

nova (mag 16) on CCD images obtained on June 21 and 22 with the Orange County Observers’ 0.56-m telescope. SN 1998co is located 3′′ east and 5′′ north of nucleus of NGC 7131 (α = 21^{h}47^{m}.7, δ = −13^{o}11′, equinox 2000.0).

P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 1998co, obtained by M. Calkins with the Fred L. Whipple Observatory (FLWO) 1.5-m telescope on June 23.4 UT, shows strong Si II absorption with a minimum at 623.63 nm and is clearly that of a type-Ia supernova near maximum. The NASA Extragalactic Database gives a redshift for NGC 7131 of 5458 km/s, so that the atmospheric expansion velocity is 10 300 km/s. CCD imaging from the FLWO 1.2-m telescope gives an offset of the supernova from the galaxy nucleus of 2′′.4 east and 4′′.6 north.

SUPERNOVA 1998cn IN NGC 3735 M. Schwartz corrects his offset figures given on IAUC 6944 to ∼ 10′′

west and 6′′ north of the galaxy’s nucleus.

SGR 1627–41 D. A. Smith and A. M. Levine report on behalf of the RXTE/ASM

teams at Massachusetts Institute of Technology and Goddard Space Flight Center: “Three short (< ∼ 1 s) burst events, most likely produced by the new soft γ-ray repeater SGR 1627–41 (IAUC 6944), were detected in SSC 3 of the RXTE/ASM at May 17.943917, 17.954243, and 18.010803 UT. These bursts are not associated with the times of any reported BATSE triggers. Due to low signal-to-noise, our standard blind-search analysis procedure did not yield a position of the source, so we performed a search, assuming that the source is located within a 1^{o}-long segment of the IPN annulus (Hurley et al., IAUC 6948) centered on the supernova remnant G337.0–0.1 noted by Woods et al. on IAUC 6948. We find a maximum reduction in χ^{2} for sources located at α = 16^{h}35^{m}.8, δ = −47^{o}32′ for event 1 and α = 16^{h}35^{m}.8, δ = −47^{o}39′ for event 2 (1σ error of 4′ in δ in both cases). Both positions are consistent with the location of G337.0–0.1. If we assume that these two bursts came from G337.0–0.1, their fluences are 2 and 3 × 10^{−7} erg/cm^{2} (5–12 keV). No flux was detected in the band 1.5–5 keV with 2σ upper limits of 1 × 10^{−7} erg/cm^{2} for both events.”

1998 June 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green